dx/(z^2+2y)=dy/(z^2+2x)=dz/-z


1
Expert's answer
2021-11-02T18:47:30-0400

Solution

"\frac{dx}{ (z^2+2y)}=\frac{dy}{ (z^2+2x)}=\frac{dz}{ (-z)}"

From first two

"(z^2+2x)dx=(z^2+2y)dy"

Here z is acting as constant , further integrating both sides we get

"z^2x+x^2+c_1=z^2y+y^2+c_2" ......(1)


Now taking first and last

"\frac{dx}{ (z^2+2y)}=\frac{dz}{ (-z)}"

Integrating both sides, here y and z will act as constant

We get

"\frac{x}{z^2+2y}+c_3=-ln(z)+c_4" ..........(2)


Equation (1) and (2) are answers .